# QUEST-A: Untrained Filtering with Trained Focusing led to Enhanced Quantum Architectures

### v1

@@ 3.3 Focused Architecture Optimization
Within the evolutionary algorithm framework of QUEST-A, parameter reuse serves not merely as a means to accelerate training but as an effective mechanism for propagating the advantages of each generation of QAS. Specifically, in each generation of candidate circuits, the best-performing individual is designated as the training agent. After undergoing intensive optimization, its parameters are selectively reused by other individuals in the next generation, thus disseminating superior features throughout the entire population.
@@
where $E(c)$ is the fitness function incorporating both the DAG path count $P(c)$ and KL divergence $D(c)$:
$E(c)=f(P(c),D(c)).$ (6)
@@
$\theta_{i}^{new}=\alpha\theta_{best}+(1-\alpha)\theta_{i}.$ (7)
@@
The rationale behind this method is twofold: firstly, based on the characteristics of evolutionary algorithms, individuals within the same generation typically possess similar circuit structures, providing a natural foundation for parameter reuse. Secondly, the re-uploading circuit template adopted by QUEST-A enables almost lossless parameter transfer between structurally similar circuits. This method achieves efficient transmission of advantageous features at the parameter level, not only improving search efficiency but also enhancing the algorithm’s ability to explore global optima in complex search spaces.
3.4 Evolutionary Synthesis of Search Strategies
A key process of the QUEST-A framework lies in its integration of the rapid architecture pruning described in Section  3.2. and the focused architecture optimization introduced in Section  3.3 into a unified evolutionary algorithm framework. This integration adopts the principles of evolutionary algorithms to guide the QAS process, combining the discrete optimization problem of circuit structures with the continuous optimization of circuit parameters.
In this two-stage joint optimization process, rapid architecture pruning and focused architecture optimization work collaboratively, forming an efficient search mechanism. Rapid architecture pruning acts as a preliminary selection mechanism, identifying promising candidate architectures. Meanwhile, focused architecture optimization conducts a refined adaptation process, fully exploiting the potential of each candidate architecture through parameter optimization. The alternation of these two processes creates a natural and efficient evolutionary cycle, driving the entire search process to continuously evolve towards superior solutions.
In the evolutionary framework of QUEST-A, the quantum circuit structure and parameters collectively constitute the genes of individuals, providing a rich space for evolutionary operations. Rapid architecture pruning and focused architecture optimization form the environment for individual survival, with DAG path count, KL divergence, and loss function values serving as multi-level fitness evaluation indicators.
This framework achieves a combination of discrete structure search and continuous parameter optimization. Rapid architecture pruning resembles large-scale structural mutation, while focused architecture optimization corresponds to small-scale parameter fine-tuning, ensuring both the breadth of the search space and the precision of optimization. The population characteristic of the evolutionary algorithm allows QUEST-A to simultaneously maintain and evaluate multiple candidate solutions. The supercircuit structure in QUEST-A can be viewed as a dynamically evolving gene pool, realizing effective knowledge transfer between generations through the parameter reuse mechanism, accelerating the optimization process and enhancing the algorithm’s adaptability.
The adaptability of the evolutionary algorithm enables QUEST-A to dynamically adjust strategies during the search process, gradually focusing on more promising search areas while maintaining the ability to explore global optimal solutions, effectively mitigating the common problem of premature convergence in QAS. The framework of QUEST-A is described as follows:
@@ 3.4 Evolutionary Synthesis of Search Strategies
2:  Define $E(c)=f(P(c),D(c))$Fitness function}
@@
9:     $\mathcal{P}^{\prime}_{g}\leftarrow$ Select top $K$ candidates based on $E(c)$
10:     $c_{best}\leftarrow\arg\max_{c\in\mathcal{P}^{\prime}_{g}}E(c)$
@@
12:        $\theta_{i}^{*}\leftarrow\arg\min_{\theta}\mathcal{L}(c_{i},\theta)$Gradient descent}
@@
19:  return  $\arg\max_{c\in\mathcal{P}_{G}}E(c)$
By leveraging both coarse-grained pruning and fine-grained optimization within an evolutionary framework, QUEST-A aims to overcome the challenges of computational complexity and premature convergence often encountered in QAS. To validate the effectiveness and efficiency of this approach, we now turn to a comprehensive experimental analysis across various quantum computing tasks.
@@ 4.1 Experiment Framework
A series of experiments were designed to validate the effectiveness and performance advantages of the QUEST-A method. These experiments include: signal representation tasks to demonstrate QUEST-A’s performance in generative tasks, particularly in handling complex signals; image classification tasks to verify QUEST-A’s performance in practical application scenarios; and variational quantum eigensolver (VQE) tasks to showcase its potential in quantum advantage domains. In each experiment, models were trained five times, with the best performance selected as the comparison result. Detailed information for experiments can be found in  B.
@@
4.2 Task I: QUEST-A in Insignal Representation
The insignal representation task aims to learn the intrinsic representation of signals through modeling, with its core objective being the learning of a function that maps input coordinates to corresponding signal values. This task can be viewed as a specialized form of continuous function approximation or signal reconstruction. In the field of quantum computing, this task demonstrates the capability of quantum models in continuous function approximation and also showcases potential in implicit neural representations tasks Mildenhall et al. (2021).
This experiment utilizes the same audio and image datasets as the benchmark models Zhao et al. (2024), ensuring consistency in data sources and preprocessing methods. Specifically, for the image representation task, images of astronaut, camera, and coffee from the scikit-image dataset Van der Walt et al. (2014) were selected.These images were first downsampled to a resolution of $32×32$ pixels and then resized through a linear layer before being input into a 10-qubit quantum model. The model processed the data through three iterations of data re-uploading for learning. Additionally, the model is equipped with two variational layers, structurally identical to the re-uploading circuits, both at the initial stage and prior to the measurement stage. These layers serve to optimize the initial state of quantum evolution and the basis vectors for measurement.
@@ 4.2 Task I: QUEST-A in Insignal Representation
The results demonstrate that in the initial stages of the QUEST-A method’s search process, the candidate circuit population obtained through rapid screening performed poorly, capturing almost no effective information. However, as iterations progressed, the performance of candidate circuits showed a rapid improvement trend. For the best-performing image, QUEST-A reduced the MSE from $4.0\times 10^{-3}$ of the baseline model to $2.2\times 10^{-3}$, achieving approximately a $45\%$ reduction in error rate. In the Astronaut image, which exhibits a more complex frequency pattern, QUEST-A similarly achieved a low error rate, demonstrating visually discernible advantages in capturing image details. Compared to the two baseline methods, QUEST-A exhibited superior performance in processing high-frequency information relative to the SOTA quantum methods, resulting in more precise retention of details and edge information in the generated images. The performance of QUEST-A stems from its further enhancement of the inherent advantages of data re-uploading quantum circuits in continuous function approximation.
@@
During the training process, we observed that the QUEST-A model quickly learned low-frequency patterns in the initial stages, similar to its performance in the image representation task. In the subsequent training process, the model made fine adjustments to the high-frequency components of the waveform and rapidly converged. This learning pattern further demonstrates QUEST-A’s adaptability in handling different types of insignal.
@@ 4.3 Task II: QUEST-A in Image Classification
Unlike signal representation tasks that focus on function fitting, image classification tasks emphasize the model’s understanding of data distribution and its ability to construct decision boundaries. This study uses the widely adopted MNIST and Fashion-MNIST datasets as benchmarks, comparing QUEST-A with several state-of-the-art quantum baseline methods, including QUILT Silver et al. (2022), Élivágar Anagolum et al. (2024), and QNAS Wang et al. (2022). To thoroughly evaluate QUEST-A’s performance in the NISQ era, we designed three experimental scenarios: numerical simulation, noise simulation, and experiments under real noisy environments. For the simulated noisy environment, we employed a composite noise model incorporating bit flip, phase flip, and depolarizing channel noise.
@@
    For baseline works, if multiple results were available under different conditions (e.g. different basic gate sets or noise environments), the best performance was selected as the final performance for the baseline model.
@@
QUEST-A demonstrated significant performance and robustness on both MNIST and Fashion-MNIST datasets. Compared to baseline models, QUEST-A achieved higher classification accuracy in noiseless environments (NNS). Notably, when noise was introduced (SNS), QUEST-A’s performance showed only a slight decrease. This noise robustness became more evident as task difficulty increased; when transitioning from 2-class to 4-class classification, QUEST-A’s performance degradation was less pronounced than that of baseline models. QUEST-A performed well on both MNIST and Fashion-MNIST datasets, indicating its generalization capability and adaptability. Further validation of QUEST-A’s performance in classification tasks is presented as Tab. 3 and Fig. 9.
@@
In-depth analysis of image multi-classification tasks further validates the robustness of the QUEST-A model. QUEST-A demonstrated high performance stability across classification tasks ranging from $2$ to $10$ classes. Although model performance showed an expected declining trend as the number of classes increased, QUEST-A still maintained a high accuracy of over $75\%$ in the most challenging 10-class classification task. A noteworthy phenomenon is that increasing the number of trainable layers did not bring about the expected significant performance improvement. On the contrary, in some cases, a slight performance decline was observed. This unexpected result may stem from two factors: firstly, the model may have experienced a certain degree of overfitting in determining decision boundaries; secondly, the current task settings may not have fully utilized the potential of the model’s assumptions.
Above experimental results not only further confirm QUEST-A’s excellent robustness in handling multi-classification tasks but also provide insights into the model’s potential performance in more complex application scenarios.
@@ 4.4 Task III: QUEST-A in Variational Quantum Eigensolver
This experiment selects three typical systems as VQE objects: the $H_{2}$ molecule (4 qubits), the Heisenberg model (5 qubits), and the Transverse Field Ising Model (TFIM, 6 qubits). Due to the unique characteristics of the VQE, QUEST does not employ data re-uploading circuits in this task. The gate set used for the task includes RX, RY, RZ, and CNOT gates. The Hamiltonian for the $H_{2}$ molecule follows the construction method from previous work Kandala et al. (2017). The Hamiltonians for TFIM and the Heisenberg model are defined as follows
@@
Heisenberg 5 $8.9\times 10^{-2}$ $1.2\times 10^{-3}$ $1.7\times 10^{-14}$
TFIM 6 $9.0\times 10^{-2}$ $1.2\times 10^{-5}$ $2.8\times 10^{-8}$
@@
QUEST-A achieved the highest precision in all test cases, demonstrating an advantage of up to 11 orders of magnitude over state-of-the-art methods. Compared to similar fast architecture search methods, QUEST-A not only successfully completed the ground state energy prediction task within chemical accuracy $(1.6\times 10^{-3}Ha)$ but also achieved more precise predictions. As the system size increased from $4$ qubits to $6$ qubits, QUEST-A maintained its performance advantage. This outcome reflects QUEST-A’s ability in preparing complex quantum states and approximating the ground state of target Hamiltonians. Its exceptional performance may stem from QUEST-A’s ability to avoid getting trapped in local optima during the search process, as well as its enhanced sensitivity to the fine structure of quantum states. Furthermore, the fact that this task did not utilize existing high-performance ansatzes corroborates that QUEST-A does not rely on specific circuit patterns, but rather can adaptively discover quantum circuit structures suitable for specific problems. This enhanced expressive capability gives QUEST-A potential advantages in handling more complex quantum systems.
In conclusion, we evaluated QUEST-A across three distinct tasks: insignal representation, image classification, and VQE. In the insignal representation task, QUEST-A achieved up to $81.8\%$ performance improvement compared to SOTA methods, with a slightly lower parameter count. For image classification, it demonstrated superior accuracy and robustness, maintaining over $75\%$ accuracy in $10$-class tasks and showing adaptability to noisy environments. In VQE tasks, QUEST-A outperformed existing methods by up to $11$ orders of magnitude in precision. These results consistently demonstrate QUEST-A’s performance and adaptability in handling complex quantum tasks, validating its effectiveness as a QAS method.
@@ 5 Discussion
The QUEST-A framework demonstrates how quantum architecture search can effectively bridge quantum and classical computing paradigms. While quantum computation excels at rapid estimation and parallel processing (as shown in KL divergence calculation), classical optimization techniques provide reliable solutions for parameter tuning and precise optimization. This complementary approach provides guiding principles for developing hybrid algorithms. Another exploration reveals that QAS can be utilized to optimize not only intermediate quantum circuits but also data encoding and measurement schemes. Recent works Wang et al. (2023); Thanos et al. (2024) in automated circuit design, combined with advances in quantum embedding Havlíček et al. (2019) and measurement Quek et al. (2021), indicate promising directions for developing task-specific quantum architectures.
Furthermore, the flexibility of the QUEST-A framework in the rapid screening phase allows for easy incorporation of hardware-specific factors. However, caution must be exercised when considering hardware noise in QAS. Although many current works Wang et al. (2022); Anagolum et al. (2024); Lu et al. (2023); He et al. (2024a) consider quantum hardware characteristics such as qubit architecture and noise properties, the time-varying nature of quantum noise may quickly render static noise models obsolete, leading to a loss of circuit adaptability. Therefore, future research needs to develop more dynamic and adaptive circuit noise mitigation strategies to better address noise challenges in practical quantum systems.

### v2

@@ 3.3 Trained Focusing
Within the evolutionary algorithm framework of QUEST-A, parameter reuse serves as an effective mechanism for knowledge transfer and accumulation across generations. Specifically, in each generation of candidate circuits, the best-performing individual is designated as the focusing agent. After fine-grained trained focusing, its parameters are selectively reused by other individuals in the next generation, thus disseminating valuable knowledge throughout the entire population.
@@
where $F(c)$ is the fitness function incorporating both the DAG path count $P(c)$ and circuits expressivity $E(c)$:
$Fitness(c)=f(P(c),E(c)).$ (6)
@@
$\theta_{i}^{new}=\alpha\theta_{best}+(1-\alpha)\theta_{i},$ (7)
@@
The rationale behind this method is twofold: firstly, based on the characteristics of evolutionary algorithms, individuals within the same generation typically possess similar circuit structures, providing a natural foundation for parameter reuse. Secondly, the re-uploading circuit template adopted by QUEST-A enables almost lossless parameter transfer between structurally similar circuits. This method achieves efficient knowledge transfer between architectures, not only improving search efficiency but also enhancing the algorithm’s ability to explore global optima in complex search spaces.
3.4 Evolutionary Synthesis of Knowledge Transfer
A key process of the QUEST-A framework lies in its integration of the untrained filtering described in Section  3.2. and trained focusing introduced in Section  3.3 into a unified evolutionary algorithm framework. This integration adopts the principles of evolutionary algorithms to guide the QAS process, combining the discrete optimization problem of circuit structures with the continuous optimization of circuit parameters.
In this knowledge-driven search process, untrained filtering and trained focusing work collaboratively, forming an efficient search mechanism. Rapid architecture pruning acts as a preliminary selection mechanism, identifying promising candidate architectures. Meanwhile, focused architecture optimization conducts a refined adaptation process, fully exploiting the potential of each candidate architecture. The alternation of these two processes creates an efficient knowledge transfer cycle, driving the entire search process to continuously evolve towards superior solutions. The framework of QUEST-A is described in Algorithm  1.
@@ 3.4 Evolutionary Synthesis of Knowledge Transfer
2:  Define $Fitness(c)=f(P(c),E(c))$Fitness function}
@@
9:     $\mathcal{P}^{\prime}_{g}\leftarrow$ Select top $K$ candidates based on $Fitness(c)$
10:     $c_{best}\leftarrow\arg\max_{c\in\mathcal{P}^{\prime}_{g}}Fitness(c)$
@@
12:        $\theta_{i}^{*}\leftarrow\arg\min_{\theta}\mathcal{L}(c_{i},\theta)$Trained focusing}
@@
19:  return  $\arg\max_{c\in\mathcal{P}_{G}}Fitness(c)$
In the evolutionary framework of QUEST-A, the quantum circuit structure and parameters collectively constitute the genes of individuals, providing a rich space for evolutionary operations. Untrained filtering and trained focusing form the environment for individual survival, with DAG path count, KL divergence, and loss function values serving as multi-level fitness evaluation indicators.
This framework achieves a synergy between coarse-grained and fine-grained search processes. Untrained filtering explores broad regions of the search space, while trained focusing refines promising areas, ensuring both search breadth and precision. The population-based evolutionary process allows QUEST-A to maintain and evaluate multiple search regions simultaneously. The supercircuit structure in QUEST-A represents a dynamically evolving search space, where knowledge transfer between generations guides the exploration towards promising architecture regions, enhancing the search efficiency and adaptability. To validate the effectiveness and efficiency of this approach, we now turn to a comprehensive experimental analysis across various quantum computing tasks.
@@ 4.1 Experiment Framework
A series of experiments were designed to validate the effectiveness of QUEST-A’s knowledge-driven search strategy. These experiments include: signal representation tasks to demonstrate QUEST-A’s performance in generative tasks, particularly in handling complex signals; image classification tasks to verify QUEST-A’s performance in practical application scenarios; and variational quantum eigensolver (VQE) tasks to showcase its potential in quantum advantage domains. In each experiment, models were trained five times, with the best performance selected as the comparison result. Detailed information for experiments can be found in  B.
@@
4.2 Task I: QUEST-A in signal Representation
The signal representation task aims to learn the intrinsic representation of signals, with its core objective being the learning of a function that maps input coordinates to corresponding signal values. This task can be viewed as a specialized form of continuous function approximation or signal reconstruction. In the field of quantum computing, this task demonstrates the capability of quantum models in continuous function approximation and also showcases potential in implicit neural representations tasks Mildenhall et al. (2021).
This experiment utilizes the same audio and image datasets as the benchmark models Zhao et al. (2024), ensuring consistency in data sources and preprocessing methods. Specifically, for the image representation task, images of astronaut, camera, and coffee from the scikit-image dataset Van der Walt et al. (2014) were selected.These images were first downsampled to a resolution of $32\times 32$ pixels and then resized through a linear layer before being input into a 10-qubit quantum model. The model processed the data through three iterations of data re-uploading for architecture discovery. Additionally, the model is equipped with two variational layers, structurally identical to the re-uploading circuits, both at the initial stage and prior to the measurement stage. These layers serve to optimize the initial state of quantum evolution and the basis vectors for measurement.
@@ 4.2 Task I: QUEST-A in signal Representation
The results demonstrate that in the initial stages of the QUEST-A method’s search process, the candidate circuit population obtained through untrained filtering performed poorly, capturing almost no effective information. However, as knowledge accumulated through iterations, the performance of candidate circuits showed a rapid improvement trend. For the best-performing image, QUEST-A reduced the MSE from $4.0\times 10^{-3}$ of the baseline model to $2.2\times 10^{-3}$, achieving approximately a $45\%$ reduction in error rate. In the Astronaut image, which exhibits a more complex frequency pattern, QUEST-A similarly achieved a low error rate, demonstrating visually discernible advantages in capturing image details. Compared to the two baseline methods, QUEST-A exhibited superior performance in processing high-frequency information relative to the SOTA quantum methods, resulting in more precise retention of details and edge information in the generated images. The performance of QUEST-A stems from its further enhancement of the inherent advantages of data re-uploading quantum circuits in continuous function approximation.
@@
During the training process, we observed that the QUEST-A model quickly learned low-frequency patterns in the initial stages, similar to its performance in the image representation task. In the subsequent training process, the model made fine adjustments to the high-frequency components of the waveform and rapidly converged. This learning pattern further demonstrates QUEST-A’s adaptability in handling different types of signal.
@@ 4.3 Task II: QUEST-A in Image Classification
Unlike signal representation tasks that focus on pattern discovery, image classification tasks emphasize the search for optimal decision boundaries. This study uses the widely adopted MNIST and Fashion-MNIST datasets as benchmarks, comparing QUEST-A with several state-of-the-art quantum baseline methods, including QUILT Silver et al. (2022), Élivágar Anagolum et al. (2024), and QNAS Wang et al. (2022). To thoroughly evaluate QUEST-A’s performance in the NISQ era, we designed three experimental scenarios: numerical simulation, noise simulation, and experiments under real noisy environments. For the simulated noisy environment, we employed a composite noise model incorporating bit flip, phase flip, and depolarizing channel noise.
@@
    For baseline works, if multiple results were available under different conditions (e.g. different basic gate sets), the best performance was selected as the final performance for the baseline model.
@@
QUEST-A demonstrated superior search capability and robustness on both MNIST and Fashion-MNIST datasets. Compared to baseline models, QUEST-A achieved higher classification accuracy in noiseless environments (NNS). Notably, when noise was introduced (SNS), QUEST-A’s performance showed only a slight decrease. This noise robustness became more evident as task difficulty increased; when transitioning from 2-class to 4-class classification, QUEST-A’s performance degradation was less pronounced than that of baseline models. QUEST-A performed well on both MNIST and Fashion-MNIST datasets, indicating its generalization capability and adaptability. Further validation of QUEST-A’s performance in classification tasks is presented as Tab. 3 and Fig. 9.
@@
In-depth analysis of image multi-classification tasks further validates the robustness of the QUEST-A model. QUEST-A demonstrated high performance stability across classification tasks ranging from $2$ to $10$ classes. Although model performance showed an expected declining trend as the number of classes increased, QUEST-A still maintained a high accuracy of over $75\%$ in the most challenging 10-class classification task. A noteworthy phenomenon is that increasing the number of trainable layers did not bring about the expected significant performance improvement. On the contrary, in some cases, a slight performance decline was observed. This unexpected result suggests that the model may have experienced a certain degree of overfitting in determining decision boundaries.
Above experimental results not only further confirm QUEST-A’s robustness in handling multi-classification tasks but also provide insights into the model’s potential performance in more complex application scenarios.
@@ 4.4 Task III: QUEST-A in Variational Quantum Eigensolver
This experiment selects three typical systems as VQE objects: the $H_{2}$ molecule (4 qubits), the Heisenberg model (5 qubits), and the Transverse Field Ising Model (TFIM, 6 qubits). Due to the unique characteristics of the VQE, QUEST does not employ data re-uploading circuits in this task. The gate set used for the task includes $RX$, $RY$, $RZ$, and $CNOT$ gates. The Hamiltonian for the $H_{2}$ molecule follows the construction method from previous work Kandala et al. (2017). The Hamiltonians for TFIM and the Heisenberg model are defined as follows:
@@
Heisenberg 5 $8.9\times 10^{-2}$ $1.2\times 10^{-3}$ $\mathbf{1.7\times 10^{-14}}$
TFIM 6 $9.0\times 10^{-2}$ $1.2\times 10^{-5}$ $\mathbf{2.8\times 10^{-8}}$
@@
    The model consistently failed to complete the VQE task for $H_{2}$ in experimental reproduction.
  • b
@@
QUEST-A achieved the highest precision in all test cases, demonstrating an advantage of up to 11 orders of magnitude over state-of-the-art methods. Compared to similar fast architecture search methods, QUEST-A not only successfully completed the ground state energy prediction task within chemical accuracy $(1.6\times 10^{-3}Ha)$ but also achieved more precise predictions. As the system size increased from $4$ qubits to $6$ qubits, QUEST-A maintained its performance advantage. This outcome reflects QUEST-A’s effectiveness in searching complex quantum architectures for ground state approximation. Its exceptional performance may stem from QUEST-A’s ability to avoid getting trapped in local optima during the search process, as well as its enhanced sensitivity to the fine structure of quantum states.
Furthermore, the fact that this task did not utilize existing high-performance ansatzes corroborates that QUEST-A’s knowledge-driven search strategy is template-independent, but rather can adaptively discover quantum circuit structures suitable for specific problems.
In conclusion, we evaluated QUEST-A’s knowledge-driven search strategy across three distinct tasks: insignal representation, image classification, and VQE. In the insignal representation task, QUEST-A achieved up to $81.8\%$ performance improvement compared to SOTA methods, with a simpler architecture space. For image classification, it demonstrated superior search capability and robustness, maintaining over $75\%$ accuracy in $10$-class tasks and showing adaptability to noisy environments. In VQE tasks, QUEST-A outperformed existing methods by up to $11$ orders of magnitude in precision. These results consistently demonstrate the effectiveness of QUEST-A’s progressive knowledge learning approach in quantum architecture search.
@@ 5 Discussion
The QUEST-A framework demonstrates how knowledge-driven search can effectively bridge quantum and classical computing paradigms. While quantum computation excels at rapid estimation and parallel processing, classical optimization techniques provide reliable solutions for parameter tuning and precise optimization. This complementary approach provides guiding principles for developing hybrid algorithms. Another exploration reveals that our retrieval-based approach can discover optimal structures for not only intermediate quantum circuits but also data encoding and measurement schemes. Recent works Wang et al. (2023); Thanos et al. (2024) in automated circuit design, combined with advances in quantum embedding Havlíček et al. (2019) and measurement Quek et al. (2021), indicate promising directions for developing task-specific quantum architectures.
Furthermore, the flexibility of the QUEST-A framework allows for incorporation of hardware-specific factors. However, caution must be exercised when considering hardware noise in QAS. Although many current works Wang et al. (2022); Anagolum et al. (2024); Lu et al. (2023); He et al. (2024a) consider quantum hardware characteristics such as qubit architecture and noise properties, the time-varying nature of quantum noise challenges the effectiveness of static search strategies. This suggests the need for more adaptive knowledge accumulation mechanisms in noisy environments, which could be a key focus for future QAS.